# Scalable Autoregressive Image Generation with Mamba

### v4

@@ Abstract
Figure 1: Autoregressive Image Generation with Mamba. We show samples from our class-conditional AiM-XL model trained on ImageNet at 256$\times$256 resolution.
@@
[…]
@@ Introduction
$[\alpha_{i},\;\beta_{i},\;\gamma_{i}]^{T}=\text{Swish}(c)W_{i}+b_{i}\quad\in\mathbb{R}^{3\times d}$ (5)
@@
$[\alpha,\;\beta,\;\gamma]^{T}=\text{Swish}(c)W+b\quad\in\mathbb{R}^{3\times d}$ (7)
@@
$[\alpha_{i},\;\beta_{i},\;\gamma_{i}]^{T}=\text{Swish}(c)W+b_{i}\quad\in\mathbb{R}^{3\times d}$ (8)
@@
$[\alpha_{i},\;\beta_{i},\;\gamma_{i}]^{T}=\text{Swish}(c)W_{j}+b_{i}\quad\in\mathbb{R}^{3\times d}$ (9)
@@
where $N$ is the total number of tokens. This approach ensures that the model effectively learns to predict each token in the sequence based on the previous tokens and class label.
@@
AiM-B Params. 93M 134M $+$44.1%
FID $\downarrow$ 4.21 3.52 --16.4%
AiM-L Params. 322M 470M $+$46.0%
FID $\downarrow$ 2.95 2.89 --2.0%
@@
Figure 8: Inference time on ImageNet1K 256$\times$256 benchmark. Result with a batch size of 16 on the A100 GPU.

### v5

@@ Abstract
Figure 1: Autoregressive Image Generation with Mamba. We show samples from our class-conditional AiM-XL model trained on ImageNet at $\times256$ resolution.
@@
Figure 2: AR image generation pipeline. Stage 1: Training the image tokenizer (encoder and quantizer) and decoder via image reconstruction. Stage 2: Training the AR model through causal sequence modeling. The symbol $\langle\text{C}\rangle$ represents the class embedding. Inference: Generating image tokens autoregressively by predicting the next token, which the decoder then converts into a synthesized image. The lock icon: Frozen weights.
@@ Introduction
$\alpha_{i},\;\beta_{i},\;\gamma_{i}]^{T}=\text{Swish}(c)W_{i}+b_{i}\quad\in\mathbb{R}^{3\times d}$ (5)
@@
$\alpha,\;\beta,\;\gamma]^{T}=\text{Swish}(c)W+b\quad\in\mathbb{R}^{3\times d}$ (7)
@@
$\alpha_{i},\;\beta_{i},\;\gamma_{i}]^{T}=\text{Swish}(c)W+b_{i}\quad\in\mathbb{R}^{3\times d}$ (8)
@@
$\alpha_{i},\;\beta_{i},\;\gamma_{i}]^{T}=\text{Swish}(c)W_{j}+b_{i}\quad\in\mathbb{R}^{3\times d}$ (9)
@@
where NN is the total number of tokens. This approach ensures that the model effectively learns to predict each token in the sequence based on the previous tokens and class label.
@@
AiM-B Params. 93M 134M ++44.1%
FID $\downarrow$ 4.21 3.52 -16.4%
AiM-L Params. 322M 470M ++46.0%
FID $\downarrow$ 2.95 2.89 -2.0%
@@
Figure 8: Inference time on ImageNet1K $\times256$ benchmark. Result with a batch size of 16 on the A100 GPU.